# An analysis of Wigner’s friend in the framework of quantum mechanics based on the principle of typicality

### v1

@@ 7.3 Postulates for determining the states of apparatuses in early stages
In this section, we have so far made an argument based on Postulate 3, the unitary time-evolution, and not based on Postulate 5, the principle of typicality. The application of Postulate 5 results in the non-unitary time-evolution in the measurement process, as we saw in Section 6. In the rest of this paper, we regard the states $|\Psi^{1}_{\mathrm{Total}}\rangle,|\Psi^{2}_{\mathrm{Total}}\rangle,\dots,|\Psi^{n}_{\mathrm{Total}}\rangle$ of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurements $\mathcal{M}^{1},\mathcal{M}^{2},\dots,\mathcal{M}^{n}$, respectively, as virtual states which provide a useful means for the calculation in our framework based on the principle of typicality, whereas the initial state $|\Psi^{0}_{\mathrm{init}}\rangle\otimes|\Phi^{1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle$ of the total system $\mathcal{S}_{\mathrm{Total}}$ is treated as a real state. Actually, the state $|\Psi^{n}_{\mathrm{Total}}\rangle$ of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the whole measurement has been used as a virtual state in our framework based on the principle of typicality, developed in Section 6. However, suggested by Proposition 7.5, we propose the following postulate for real states.
@@
    Let $i\in\{1,\dots,n\}$. For every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$, if the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the states of the system $\mathcal{S}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{i}$, then the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately before the measurement $\mathcal{M}^{i}$.
@@
    Let $i,j\in\{1,\dots,n\}$ with $i<j$. For every $m_{i}\in\Omega_{i}$, if the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the final states of the apparatus $\mathcal{A}_{i}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{j}$, then the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately before the measurement $\mathcal{M}^{j}$. ∎
@@
Suppose that the measurement operators $\{M^{i}_{m_{i}}\}_{m_{i}\in\Omega_{i}}$ form a PVM for every $i=1,\dots,n$. Suppose that the state of the total system $\mathcal{S}_{\mathrm{Total}}$, consisting of the system $\mathcal{S}$ and the apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$, immediately after the measurement $\mathcal{M}^{n}$ is
@@
where $m_{1}\in\Omega_{1},\dots,m_{n}\in\Omega_{n}$ and $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$ is a state of the system $\mathcal{S}$. Let $A$ be the set of $i\in\{1,\dots,n-1\}$ such that the final states of the apparatus $\mathcal{A}_{i}$ is unchanged after the measurement by the apparatus $\mathcal{A}_{n}$. We define a finite sequence $i_{1},i_{2},\dots,i_{k}\in\{1,2,\dots,n-1\}$ by the condition that $A=\{i_{1},i_{2},\dots,i_{k}\}$ and $i_{1}<i_{2}<\dots<i_{k}$. Then the state of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{n-1}$ is given by
@@
up to the normalization factor. Note that this state is just $|\Psi^{n-1}_{\mathrm{Total}}\rangle$ in the case of $A=\emptyset$, i.e., in the case of $k=0$. ∎
@@
where $\{L^{i}_{m_{i}}\}_{m_{i}\in\Omega_{i}}$ are measurement operators acting on $\mathcal{H}_{\mathcal{S}}$. It is then easy to see that, for every $i=1,\dots,n-1$, the final states of the apparatus $\mathcal{A}_{i}$ is confirmed before the measurement by the apparatus $\mathcal{A}_{i+1}$. Thus, for every $i=1,\dots,n-1$, the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ is the apparatus $\mathcal{A}_{i}$ itself.
@@
where $m_{1}\in\Omega_{1},\dots,m_{n}\in\Omega_{n}$ and $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$ is a state of the system $\mathcal{S}$. Then, according to Postulate 6 we have that, for every $i=1,\dots,n-1$, the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately before the measurement $\mathcal{M}^{i+1}$, that is, the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{i}$. This result is comprehensively used in our former works implicitly to make the operational refinements of quantum information theory in our framework based on the principle of typicality [46, 47, 49, 50].
Suppose further that the measurement operators $\{M^{i}_{m_{i}}\}_{m_{i}\in\Omega_{i}}$ form a PVM for every $i=1,\dots,n$. Then, since the final states of the apparatus $\mathcal{A}_{i}$ is unchanged after the measurement by the apparatus $\mathcal{A}_{n}$ for every $i=1,\dots,n-1$, according to Postulate 7 we have that the state of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{n-1}$ is given by
@@
up to the normalization factor. This result is also comprehensively used in our former works implicitly to make the operational refinements of quantum information theory in our framework based on the principle of typicality [46, 47, 49, 50]. In particular, the result (22) implicitly played a crucial role in the derivations of Postulate 8 of Tadaki [47], which is a refined rule of the Born rule for mixed states, from the principle of typicality in several scenarios of the setting of measurements, as described in Sections 11 and 12 of Tadaki [47]. Thus, this effectiveness of the result (22) supports general use of Postulate 7 in a specific situation such as Wigner’s friend and Deutsch’s thought experiment.
@@ 8 Wigner’s friend
After the measurement $\mathcal{M}^{\mathcal{F}}$ by $\mathcal{F}$ over the system $\mathcal{S}$, the observer $\mathcal{W}$ performs a measurement $\mathcal{M}^{\mathcal{W}}$ described by the measurement operators $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$ over the system $\mathcal{F}$ (according to Postulate 4), where
@@
Applying Postulate 4 to the measurement $\mathcal{M}^{\mathcal{F}}$, immediately after the measurement $\mathcal{M}^{\mathcal{F}}$, either the following two possibilities (i) or (ii) occurs:
@@
According to (24), immediately after the measurement $\mathcal{M}^{\mathcal{F}}$, the state of the composite system $\mathcal{S}+\mathcal{F}$ consisting of $\mathcal{S}$ and $\mathcal{F}$ is
$\frac{1}{\sqrt{2}}(|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle+|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle).$ (26)
@@ 9 Analysis of Wigner’s friend based on the principle of typicality
where $I_{\mathcal{S}}$ denotes the identity operator acting on $\mathcal{H}_{\mathrm{S}}$. Thus, the finite collection $\{M_{k,l}\}_{(k,l)\in\Omega}$ forms measurement operators. We denote the identity operator acting on $\mathcal{H}_{\mathrm{W}}$ by $I_{\mathcal{W}}$. Then the sequential applications of $U_{\mathrm{I}}$ and $U_{\mathrm{II}}$ result in:
@@ 9.1 Application of the principle of typicality
Now, let us apply Postulate 5, the principle of typicality, to the setting of measurements developed above. Let $\omega$ be our world in the infinite repetition of the measurement described by the measurement operators $\{M_{k,l}\}_{(k,l)\in\Omega}$ in the above setting. This $\omega$ is an infinite sequence over $\Omega$ consisting of records in the two observers $\mathcal{F}$ and $\mathcal{W}$ which is being generated by the infinite repetition of the measurements described by the measurement operators $\{M_{k,l}\}_{(k,l)\in\Omega}$ in the above setting. Since the Bernoulli measure $\lambda_{P}$ on $\Omega^{\infty}$ is the probability measure induced by the probability measure representation for the prefixes of worlds in the above setting, it follows from Postulate 5 that $\omega$ is Martin-Löf $P$-random.
We use $\alpha$ and $\beta$ to denote the infinite sequences over $\{0,1\}$ and $\{0,1,2\}$, respectively, such that $(\alpha(n),\beta(n))=\omega(n)$ for every $n\in\mathbb{N}^{+}$. Then it follows from (30) and Corollary 4.4 that
@@
Let us determine the states of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in our world $\omega$. Note first that the condition (10) holds in our setting of measurements developed above. On the one hand, according to Definition 7.1, it is easy to check that the states of the system $\mathcal{S}$ is unchanged after the measurement by the observer $\mathcal{W}$ and therefore the states of the system $\mathcal{S}$ is confirmed before the measurement by the observer $\mathcal{W}$, where “apparatus” in Definition 7.1 should read “observer”. Using (32), we see that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the system $\mathcal{S}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is $|\alpha(n)\rangle$ in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$. Thus, it follows from Postulate 6 (i) that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is
@@
On the other hand, according to Definition 7.3, it is easy to check that the final states of the observer $\mathcal{F}$ is unchanged after the measurement by the observer $\mathcal{W}$ and therefore the final states of the observer $\mathcal{F}$ is confirmed before the measurement by the observer $\mathcal{W}$, where “apparatus” in Definition 7.3 should read “observer”. Using (32), we see that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is $|\Phi^{\mathcal{F}}[\alpha(n)]\rangle$ in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$. Thus, it follows from Postulate 6 (ii) that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is
@@
In addition, it follows from (30) and Theorem 4.1 that $\alpha$ is a Martin-Löf $U$-random infinite binary sequence, where $U$ is a finite probability space on $\{0,1\}$ such that $U(0)=U(1)=1/2$. Namely, $\alpha$ is a Martin-Löf random infinite binary sequence. Therefore, it follows from Theorem 4.2 that for every $k\in\{0,1\}$ it holds that
@@
where $N_{k}(\sigma)$ denotes the number of the occurrences of $k$ in $\sigma$ for every $k\in\{0,1\}$ and every $\sigma\in\{0,1\}^{*}$, as expected from the point of view of the conventional quantum mechanics. Thus, in our world $\omega$, for each $k=0,1$ the following (i) and (ii) simultaneously occur in half of the infinite repetitions of the experiment: (i) The observer $\mathcal{F}$ records the value $k$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$, and (ii) the observer $\mathcal{F}$ records the value $k$ and the observer $\mathcal{W}$ knows that the observer $\mathcal{F}$ records the value $k$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$, in accordance with our common sense.
We can use Postulate 7 to determine the states of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in our world $\omega$, instead of using Postulate 6. First note that the measurement operators $\{M^{\mathcal{F}}_{0},M^{\mathcal{F}}_{1}\}$ and $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$ form PVMs. On the one hand, applying $U_{\mathrm{I}}\otimes I_{\mathcal{W}}$ to the initial state $|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle$ of the repeated once of the experiments results in a state $|\Psi^{\mathcal{F}}_{\mathrm{Total}}\rangle$ given by
@@
where the second and last equalities follow from (24) and (23), respectively. On the other hand, the final states of the observer $\mathcal{F}$ is unchanged after the measurement by the observer $\mathcal{W}$, as we saw above. Thus, it follows from Postulate 7 and (32) that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
@@
$P\left(|\Phi^{\mathcal{F}}[\alpha(n)]\rangle\right):=I_{\mathcal{S}}\otimes\lvert\Phi^{\mathcal{F}}[\alpha(n)]\rangle\langle\Phi^{\mathcal{F}}[\alpha(n)]\rvert\otimes I_{\mathcal{W}}.$

### v2

@@ 7.3 Postulates for determining the states of apparatuses in early stages
In the preceding subsection, we have made an argument based on Postulate 3, the unitary time-evolution, and not based on Postulate 5, the principle of typicality. The application of Postulate 5 results in the non-unitary time-evolution in the measurement process, as we saw in Section 6. In the rest of this paper, we regard the states $|\Psi^{1}_{\mathrm{Total}}\rangle,|\Psi^{2}_{\mathrm{Total}}\rangle,\dots,|\Psi^{n}_{\mathrm{Total}}\rangle$ of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurements $\mathcal{M}^{1},\mathcal{M}^{2},\dots,\mathcal{M}^{n}$, respectively, as virtual states which provide a useful means for the calculation in our framework based on the principle of typicality, whereas the initial state $|\Psi^{0}_{\mathrm{init}}\rangle\otimes|\Phi^{1}_{\mathrm{init}}\rangle\otimes\dots\otimes|\Phi^{n}_{\mathrm{init}}\rangle$ of the total system $\mathcal{S}_{\mathrm{Total}}$ is treated as a real state. Actually, the state $|\Psi^{n}_{\mathrm{Total}}\rangle$ of the total system $\mathcal{S}_{\mathrm{Total}}$ at the instant when the whole measurement consisting of $\mathcal{M}^{1},\mathcal{M}^{2},\dots,\mathcal{M}^{n}$ is completed has been regarded as a virtual state in our framework based on the principle of typicality, developed in Section 6. However, suggested by Proposition 7.5, we propose the following postulate for real states.
@@
    Let $i\in\{1,\dots,n\}$. For every $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$, if the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the states of the system $\mathcal{S}$ are confirmed before the measurement by the apparatus $\mathcal{A}_{i}$, then the state of the system $\mathcal{S}$ is $|\Psi\rangle$ immediately before the measurement $\mathcal{M}^{i}$.
@@
    Let $i,j\in\{1,\dots,n\}$ with $i<j$. For every $m_{i}\in\Omega_{i}$, if the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{n}$ and the final states of the apparatus $\mathcal{A}_{i}$ are confirmed before the measurement by the apparatus $\mathcal{A}_{j}$, then the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately before the measurement $\mathcal{M}^{j}$. ∎
@@
Suppose that the measurement operators $\{M^{i}_{m_{i}}\}_{m_{i}\in\Omega_{i}}$ form a PVM in $\mathcal{H}_{\mathcal{S}}\otimes\overline{\mathcal{H}}_{1}\otimes\dots\otimes\overline{\mathcal{H}}_{i-1}$ for every $i=1,\dots,n$. Suppose that the state of the total system $\mathcal{S}_{\mathrm{Total}}$, consisting of the system $\mathcal{S}$ and the apparatuses $\mathcal{A}_{1},\dots,\mathcal{A}_{n}$, immediately after the measurement $\mathcal{M}^{n}$ is
@@
where $m_{1}\in\Omega_{1},\dots,m_{n}\in\Omega_{n}$ and $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$ is a state of the system $\mathcal{S}$. Let $D$ be the set of $i\in\{1,\dots,n-1\}$ such that the final states of the apparatus $\mathcal{A}_{i}$ are unchanged after the measurement by the apparatus $\mathcal{A}_{n}$. We define a finite sequence $i_{1},i_{2},\dots,i_{k}\in\{1,2,\dots,n-1\}$ by the condition that $D=\{i_{1},i_{2},\dots,i_{k}\}$ and $i_{1}<i_{2}<\dots<i_{k}$. Then the state of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{n-1}$ is given by
@@
up to the normalization factor. Note that this state is just $|\Psi^{n-1}_{\mathrm{Total}}\rangle$ in the case of $D=\emptyset$, i.e., in the case of $k=0$. ∎
@@
where $\{L^{i}_{m_{i}}\}_{m_{i}\in\Omega_{i}}$ are measurement operators acting on $\mathcal{H}_{\mathcal{S}}$. It is then easy to see that, for every $i=1,\dots,n-1$, the final states of the apparatus $\mathcal{A}_{i}$ are confirmed before the measurement by the apparatus $\mathcal{A}_{i+1}$. Thus, for every $i=1,\dots,n-1$, the confirming point for the final states of the apparatus $\mathcal{A}_{i}$ is the apparatus $\mathcal{A}_{i}$ itself.
@@
where $m_{1}\in\Omega_{1},\dots,m_{n}\in\Omega_{n}$ and $|\Psi\rangle\in\mathcal{H}_{\mathcal{S}}$ is a state of the system $\mathcal{S}$. Then, according to Postulate 6 we have that, for every $i=1,\dots,n-1$, the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately before the measurement $\mathcal{M}^{i+1}$, that is, the state of the apparatus $\mathcal{A}_{i}$ is $|\Phi^{i}[m_{i}]\rangle$ immediately after the measurement $\mathcal{M}^{i}$. This result is comprehensively used in our former works implicitly in order to make the operational refinements of quantum mechanics and quantum information theory in our framework based on the principle of typicality [46, 47, 50].
Suppose further that the measurement operators $\{M^{i}_{m_{i}}\}_{m_{i}\in\Omega_{i}}$ form a PVM in $\mathcal{H}_{\mathcal{S}}\otimes\overline{\mathcal{H}}_{1}\otimes\dots\otimes\overline{\mathcal{H}}_{i-1}$ for every $i=1,\dots,n$. Then, since the final states of the apparatus $\mathcal{A}_{i}$ are unchanged after the measurement by the apparatus $\mathcal{A}_{n}$ for every $i=1,\dots,n-1$, according to Postulate 7 we have that the state of the total system $\mathcal{S}_{\mathrm{Total}}$ immediately after the measurement $\mathcal{M}^{n-1}$ is given by
@@
up to the normalization factor. This result is also comprehensively used in our former works implicitly in order to make the operational refinements of quantum mechanics and quantum information theory in our framework based on the principle of typicality [46, 47, 50]. In particular, the result (22) implicitly played a crucial role in the derivations of Postulate 8 of Tadaki [47], which is a refined rule of the Born rule for mixed states, from the principle of typicality in several scenarios of the setting of measurements, as described in Sections 11 and 12 of Tadaki [47]. Thus, this effectiveness of the result (22) supports general use of Postulate 7 in a specific situation such as Wigner’s friend and Deutsch’s thought experiment.
@@ 8 Wigner’s friend
After the measurement $\mathcal{M}^{\mathcal{F}}$ performed by $\mathcal{F}$ over the system $\mathcal{S}$, the observer $\mathcal{W}$ performs a measurement $\mathcal{M}^{\mathcal{W}}$ described by the measurement operators $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$ over the system $\mathcal{F}$ (according to Postulate 4), where
@@
Applying Postulate 4 to the measurement $\mathcal{M}^{\mathcal{F}}$, either the following two possibilities (i) or (ii) occurs immediately after the measurement $\mathcal{M}^{\mathcal{F}}$:
@@
According to (24), the state of the composite system $\mathcal{S}+\mathcal{F}$ consisting of $\mathcal{S}$ and $\mathcal{F}$ is
$\frac{1}{\sqrt{2}}(|0\rangle\otimes|\Phi^{\mathcal{F}}[0]\rangle+|1\rangle\otimes|\Phi^{\mathcal{F}}[1]\rangle)$ (26)
immediately after the measurement $\mathcal{M}^{\mathcal{F}}$.
@@ 9 Analysis of Wigner’s friend based on the principle of typicality
where $I_{\mathcal{S}}$ denotes the identity operator acting on $\mathcal{H}_{\mathrm{S}}$. Thus, the finite collection $\{M_{k,l}\}_{(k,l)\in\Omega}$ satisfies the completeness equation, and therefore forms measurement operators. We denote the identity operator acting on $\mathcal{H}_{\mathrm{W}}$ by $I_{\mathcal{W}}$. Then the sequential applications of $U_{\mathrm{I}}$ and $U_{\mathrm{II}}$ result in:
@@ 9.1 Application of the principle of typicality
Now, let us apply Postulate 5, the principle of typicality, to the setting of measurements developed above. Let $\omega$ be our world in the infinite repetition of the measurement described by the measurement operators $\{M_{k,l}\}_{(k,l)\in\Omega}$ in the above setting. This $\omega$ is an infinite sequence over $\Omega$ consisting of records in the two observers $\mathcal{F}$ and $\mathcal{W}$ which is being generated by the infinite repetition of the measurement described by the measurement operators $\{M_{k,l}\}_{(k,l)\in\Omega}$ in the above setting. Since the Bernoulli measure $\lambda_{P}$ on $\Omega^{\infty}$ is the probability measure induced by the probability measure representation for the prefixes of worlds in the above setting, it follows from Postulate 5 that $\omega$ is Martin-Löf $P$-random.
We use $\alpha$ and $\beta$ to denote the infinite sequences over $\{0,1\}$ and $\{0,1,2\}$, respectively, such that $(\alpha(n),\beta(n))=\omega(n)$ for every $n\in\mathbb{N}^{+}$. Then, since $\omega$ is Martin-Löf $P$-random, it follows from (30) and Corollary 4.4 that
@@
Let us determine the states of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in our world $\omega$. Note first that the condition (10) holds in our setting of measurements developed above. On the one hand, according to Definition 7.1, it is easy to check that the states of the system $\mathcal{S}$ are unchanged after the measurement by the observer $\mathcal{W}$ and therefore the states of the system $\mathcal{S}$ are confirmed before the measurement by the observer $\mathcal{W}$, where “apparatus” in Definition 7.1 should read “observer”. Using (32), we see that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the system $\mathcal{S}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is $|\alpha(n)\rangle$ in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$. Thus, it follows from Postulate 6 (i) that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the system $\mathcal{S}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is
@@
On the other hand, according to Definition 7.3, it is easy to check that the final states of the observer $\mathcal{F}$ are unchanged after the measurement by the observer $\mathcal{W}$ and therefore the final states of the observer $\mathcal{F}$ are confirmed before the measurement by the observer $\mathcal{W}$, where “apparatus” in Definition 7.3 should read “observer”. Using (32), we see that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$ is $|\Phi^{\mathcal{F}}[\alpha(n)]\rangle$ in the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$. Thus, it follows from Postulate 6 (ii) that, in our world $\omega$, for every $n\in\mathbb{N}^{+}$ the state of the observer $\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is
@@
In addition, since $\omega$ is Martin-Löf $P$-random, it follows from Theorem 4.1 and (30) that $\alpha$ is a Martin-Löf $U$-random infinite binary sequence, where $U$ is a finite probability space on $\{0,1\}$ such that $U(0)=U(1)=1/2$. Namely, $\alpha$ is a Martin-Löf random infinite binary sequence. Therefore, it follows from Theorem 4.2 that for every $k\in\{0,1\}$ it holds that
@@
where $N_{k}(\alpha\!\!\restriction_{n})$ denotes the number of the occurrences of $k$ in the prefix of $\alpha$ of length $n$, as expected from the point of view of the conventional quantum mechanics. Thus, in our world $\omega$, for each $k=0,1$ the following (i) and (ii) simultaneously occur in half of the infinite repetitions of the experiment: (i) The observer $\mathcal{F}$ records the value $k$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$, and (ii) the observer $\mathcal{F}$ records the value $k$ and the observer $\mathcal{W}$ knows that the observer $\mathcal{F}$ records the value $k$ immediately after the measurement $\mathcal{M}^{\mathcal{W}}$, in accordance with our common sense.
We can use Postulate 7 to determine the states of the composite system $\mathcal{S}+\mathcal{F}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ in our world $\omega$, instead of using Postulate 6. First note that each of the measurement operators $\{M^{\mathcal{F}}_{0},M^{\mathcal{F}}_{1}\}$ and $\{M^{\mathcal{W}}_{0},M^{\mathcal{W}}_{1},M^{\mathcal{W}}_{2}\}$ forms a PVM. On the one hand, applying $U_{\mathrm{I}}\otimes I_{\mathcal{W}}$ to the initial state $|+\rangle\otimes|\Phi^{\mathcal{F}}_{\mathrm{init}}\rangle\otimes|\Phi^{\mathcal{W}}_{\mathrm{init}}\rangle$ in the repeated once of the experiments results in a state $|\Psi^{\mathcal{F}}_{\mathrm{Total}}\rangle$ given by
@@
where the second and last equalities follow from (24) and (23), respectively. On the other hand, the final states of the observer $\mathcal{F}$ are unchanged after the measurement by the observer $\mathcal{W}$, as we saw above. Thus, it follows from Postulate 7 and (32) that, in our world $\omega$, for each $n\in\mathbb{N}^{+}$ the state of the $n$th composite system $\mathcal{S}+\mathcal{F}+\mathcal{W}$ immediately after the measurement $\mathcal{M}^{\mathcal{F}}$ is given by
@@
$P\left(|\Phi^{\mathcal{F}}[\alpha(n)]\rangle\right)=I_{\mathcal{S}}\otimes\lvert\Phi^{\mathcal{F}}[\alpha(n)]\rangle\langle\Phi^{\mathcal{F}}[\alpha(n)]\rvert\otimes I_{\mathcal{W}}.$